OpenStudy (anonymous):

Which of the following represents 7x^4/5 in radical form?

OpenStudy (anonymous):

\[\sqrt[5]{7x^4}\]

OpenStudy (anonymous):

When you have a fraction as an exponent, the numerator raises the number to that power and the denominator roots it. So x^(1/2) is \[\sqrt{x ^{1}}\] Which is just \[\sqrt{x}\] x^(5/6) is x raised to the 5th power and then 6th rooted, so it's \[\sqrt[6]{x ^{5}}\]